OpenStudy (anonymous):

prove the identity 1+sin2x/sin2x=1+(1/2)secxcscx

OpenStudy (cwrw238):

take righthand side: 1+(1/2)secxcscx = 1 + 1 ------- 2 sinx cosx = 1 + 1 ---- sin 2x - this rearranges to lefthand side can you do this?

OpenStudy (cwrw238):

add them as you would simple fractions 1 + 1 -- 4 LCM = 4: = 4 + 1 ---- 4

OpenStudy (anonymous):

how do get the 4